what do you understand by ordinary paint of the equation

a0(x)d2y/dx2+a1(x)dy/dx+a2(x)y=0? hence using taylor series expansion find a series solution in powers of x for the equation(2x3-3)d2y/dx2-2xdy/dx+y=0 y(0)=-1

y1(0)=5



1
Expert's answer
2021-07-05T18:12:59-0400

Ordinary point of equation -


"a_{0}(x)\dfrac{d^{2}y}{dx^{2}}+a_1(x)\dfrac{dy}{dx}+a_2(x)y=0"



A point "x_{0}" is an ordinary point of an differential equation if both "p(x)\ and \ q(x)" are analytical at "x=0" .If point is not ordinary the it is a singular point . In other words , we can say that , a point

"x=x_{0}" is called ordinary point of equation if "a_{o}(x)\ne0" at "x=x_{0}" or other than is is singular point.

2


"=(2x^{3}-3)\dfrac{d^{2}y}{dx^{2}}-2x\dfrac{dy}{dx}+y=0"


"y(0)=-1" "y^{'}(0)=-5"


Now , we have to expand this series by taylor's series method , so we will expand it like this -


we know that general form of taylor's series , which is given by -


"=" Let "y(x)" be a solution of differential equation -


now taylor's series can be written as , we have to expand our series between point "x=0" ,


"=y(x)=y(0)+y^{'}(0)(x-0)+\dfrac{{y^{''}(0)}(x-0)^{2}}{2}" "+........."


We can write our given differential equation in this form also-


"=y^{''}(x)=\dfrac{2xy^{'}{(x)}}{(2x^{3}-3)}" "-\dfrac{y(x)}{(2x^{3}-3)}"


"=" Now putting "x=0, we\ get-"



"=y^{''}(0)=" "\dfrac{20y^{'}{(0)}}{(20^{3}-3)}-\dfrac{y(0)}{(20^{3}-3)}"


"=" "y^{''}=\dfrac{-1}{3}"



now putting all the values in general form of taylor's series ,we get -


"=" "y(x)=y(0)+y^{'}(0)(x-0)+\dfrac{{y^{''}(0)}(x-0)^{2}}{2}+......"


"=y(x)=-1+5x-\dfrac{x^{2}}{6}+......"


This is our general form of taylor.s series .